a rectangle has a perimeter of 68 inches. its legnth is 2 inches less than 3 times a width. what are the legth and width of the rectange
Answer by tran3209(100)

Step 1.
-First you must suppose that the width is "x".
-The length of the rectangle is 2 inches less than 3 times the width.
So that would make the length (3x-2)
Step 2.
-To find the perimiter of a rectangle: 2W X 2L.
Step 3.
-Form the equation.
2(3x-2) + 2(x) = 68
6x-4 + 2x = 68
8x=64
x=8
The width is 8 inches.
The length is 22 inches.
Check.
2(8)+ 2(22) = 68 inches